OpenStudy (anonymous):

(x + 1.2) + (3.75 + 2x) + (1.05 + 2x) = x + 106 HELP ME please and thank you I will give medals to who ever attempts :)

OpenStudy (anonymous):

All of it's addition at first, so just take out all of the constants and add them up. Then, do the same with the variables and you're left with 6 + 5x = x + 106. Then, subtract x from both sides to get 6 + 4x = 106. Subtract 6 from both sides to get 4x = 100, then divide by 4. You should be left with x = 25. Just isolate the variables and you will be fine in the future.